For a given set of constants α, β, and γ the functions


ƒ (x, y) =


7𝑥


𝛼𝑦


2𝑥𝛽


and ġ (x, y) =


3𝑥


𝛽−𝛾


𝑥𝛽−𝑦2


are homogeneous of degree 2 and 1 respectively. Determine the values for α, β, and γ.


1
Expert's answer
2021-11-29T17:11:25-0500

An equation of the form f(x,y)=0 is said to be the homogeneous equation of degree n, where n is a positive integer, and if for some real number k, we have f(kx,ky)=knf(x,y)


"f(x,y)=\frac{7x^{\alpha}y}{2x^{\beta}}"

"f(kx,ky)=k^2f(x,y)"

"\frac{7k^{\alpha}x^{\alpha}ky}{2k^{\beta}x^{\beta}}=\frac{7k^2x^{\alpha}y}{2x^{\beta}}"


"k^{\alpha-\beta+1}=k^2"

"\alpha-\beta=1"


"ġ (x, y) =\frac{3x^{\beta-\gamma}}{x^{\beta}-y^2}"


"\frac{3k^{\beta-\gamma}x^{\beta-\gamma}}{k^{\beta}x^{\beta}-k^2y^2}=\frac{3kx^{\beta-\gamma}}{x^{\beta}-y^2}"


"k^{\beta-\gamma}=k(k^{\beta}-k^2)"

"k^{\beta-\gamma-1}=k^{\beta}-k^2"